# Efficiently Verifiable Proofs of Data Attribution

### v1

@@ 1 Introduction
In this paper, we demonstrate an interactive two-message protocol where the third party’s computational cost, which is measured by the number of full model retrainings it must conduct, is independent of the attribution set size $N$ (e.g., in data attribution $N$ is dataset size, while in component attribution it is the number of components). That is, this number of required retrainings does not grow as a function of $N$. Our protocol comes with strong guarantees of correctness for the third party, even if the model developer deviates from the agreed protocol arbitrarily, and with infinite compute. To make this precise, we must specify (i) the form of attributions we study, (ii) how we measure computational cost, (iii) what our accuracy metric is, and finally (iv) what guarantees of correctness we can (and should) provide for our protocol.
@@
In predictive training data attribution (TDA), the goal is to answer counterfactual questions: “What would have happened to a model’s behavior if we had trained it on a different subset of the data?” To formalize this, we consider a model output function, $f:\{\pm 1\}^{N}\to\real$. The domain of $f$ are subsets $x$ of a fixed training dataset $S$ of size $N$ (where $x\in\{\pm 1\}^{N}$ and $x_{i}=1$ indicates inclusion of the $i^{th}$ data point from $S$ and $x_{i}=-1$ indicates exclusion). The model output function $f(x)$ measures the (expected) model behavior if it were trained on $S$, and is evaluated by retraining a model $\theta\sim\mathcal{T}(x)$ via a training algorithm $\mathcal{T}$, and then computing some function of $\theta$, often the prediction $\theta(z)$ on an input point $z$ (Ilyas et al.,, 2022), or the test error of $\theta$ on some subgroup of the data (Jain et al.,, 2024).
@@ 2. $\bullet$
As in the naive protocol, only checking the MSE fails to shed light on whether such cheating has been carried out by the Prover (without loss of generality on the exact cheating strategy). To emphasize the point further, consider the case that the Prover wants to advantage a subset of data contributors $A$ over all other contributors, so they receive more payment. In order to do so, they modify $\mathbf{a}$ by increasing coordinates where $i\in A$ by some value $\beta>0$ to produce a new attribution vector $a^{\prime}$. If $\beta$ is reasonably small, the modified $a^{\prime}$ might also have low MSE, and so a Verifier that merely checks the MSE is below a certain threshold would accept these attributions.
@@ 3. $\bullet$
    Efficiency. The Verifier’s workload scales independently of the dataset size $N$.
@@ 2.1 The Protocol Guarantees
The Verifier’s cost is required to be independent of $N=|S|$. For our choices of $\Phi$ and $\mathrm{err}$ (as defined above), the honest Prover’s cost can be shown to necessarily grow with $N$. Hence, our protocols allow the Verifier to expend arbitrarily less cost than the Prover.
@@ 2.2 Main Theorem
We will take the Verifier’s cost to be the expected number of training runs they conduct (over randomness of the protocol execution). [^4] This is but one way to define cost. For now, we only mention that this notion of sample complexity in data attribution is a commonly used efficiency metric in the field (see e.g, Ilyas et al., (2022); Park et al., (2023)). We emphasize this metric as each retraining constitutes a significant computational effort, potentially scaling with the dataset size $N$ itself. Hence, it captures the primary bottleneck for the Verifier. We are now ready to state the theorem.
@@ 3 A “Simple” Non-Interactive Verification Protocol
While this cost is independent of the dataset size $N$, the cubic dependence on $1/\varepsilon$ can still be substantial.
@@ 4.1 Proof Sketch of Main Theorem
Efficiency. To show efficiency, we will directly estimate the expected number of model trainings done by the Verifier, which is straightforward: it is the sum of the number of model trainings in Round $3$, which consists of $O(\log(1/\delta)/\epsilon^{2})$ spot checks, and $O(\log(1/\delta)/\epsilon^{2})$ to estimate the MSE. Thus the overall cost is $O(\log(1/\delta)/\epsilon^{2})$.
@@ A.1 Harmonic Analysis on the p-biased Cube
We consider the $p\text{-biased}$ distribution $\mathcal{B}_{p}$ over $\{\pm 1\}^{N}$, where each coordinate $x_{i}$ is independently chosen to be $1$ with probability $p$ and 1-1 with probability 1p1-p. Let $\mu=\mathbb{E}\left[x_{i}\right]=p-(1-p)=2p-1$ and $\sigma^{2}=\mathrm{Var}(x_{i})=\mathbb{E}\left[x_{i}^{2}\right]-(\mathbb{E}\left[x_{i}\right])^{2}=1-(2p-1)^{2}=4p(1-p)$. The uniform distribution corresponds to $p=1/2$, where $\mu=0$ and $\sigma^{2}=1$.
@@ Appendix B Efficient Estimation of Linear Datamodel Residual
This approach allows estimating the quality of the best linear fit using a number of samples that is proportional to $1/\varepsilon^{3}$ for a desired approximation error $\varepsilon$ but is independent of the dimension $N$.

### v2

@@ 1 Introduction
In this paper, we demonstrate an interactive two-message protocol where the third party’s computational cost, which is measured by the number of full model retrainings it must conduct, is independent of the attribution set size NN (e.g., in data attribution NN is dataset size, while in component attribution it is the number of components). That is, this number of required retrainings does not grow as a function of NN. Our protocol comes with strong guarantees of correctness for the third party, even if the model developer deviates from the agreed protocol arbitrarily, and with infinite compute. To make this precise, we must specify (i) the form of attributions we study, (ii) how we measure computational cost, (iii) what our accuracy metric is, and finally (iv) what guarantees of correctness we can (and should) provide for our protocol.
@@
In predictive training data attribution (TDA), the goal is to answer counterfactual questions: “What would have happened to a model’s behavior if we had trained it on a different subset of the data?” To formalize this, we consider a model output function, $f:\{\pm 1\}^{N}\to\real$. The domain of $f$ are subsets $x$ of a fixed training dataset $S$ of size NN (where $x\in\{\pm 1\}^{N}$ and $x_{i}=1$ indicates inclusion of the $i^{th}$ data point from $S$ and $x_{i}=-1$ indicates exclusion). The model output function $f(x)$ measures the (expected) model behavior if it were trained on $S$, and is evaluated by retraining a model $\theta\sim\mathcal{T}(x)$ via a training algorithm $\mathcal{T}$, and then computing some function of $\theta$, often the prediction $\theta(z)$ on an input point $z$ (Ilyas et al.,, 2022), or the test error of $\theta$ on some subgroup of the data (Jain et al.,, 2024).
@@ 2. $\bullet$
As in the naive protocol, only checking the MSE fails to shed light on whether such cheating has been carried out by the Prover (without loss of generality on the exact cheating strategy). To emphasize the point further, consider the case that the Prover wants to advantage a subset of data contributors AA over all other contributors, so they receive more payment. In order to do so, they modify $\mathbf{a}$ by increasing coordinates where $i\in A$ by some value $\beta>0$ to produce a new attribution vector $a^{\prime}$. If $\beta$ is reasonably small, the modified $a^{\prime}$ might also have low MSE, and so a Verifier that merely checks the MSE is below a certain threshold would accept these attributions.
@@ 3. $\bullet$
    Efficiency. The Verifier’s workload scales independently of the dataset size NN.
@@ 2.1 The Protocol Guarantees
The Verifier’s cost is required to be independent of $N=|S|$. For our choices of $\Phi$ and $\mathrm{err}$ (as defined above), the honest Prover’s cost can be shown to necessarily grow with NN. Hence, our protocols allow the Verifier to expend arbitrarily less cost than the Prover.
@@ 2.2 Main Theorem
We will take the Verifier’s cost to be the expected number of training runs they conduct (over randomness of the protocol execution). [^4] This is but one way to define cost. For now, we only mention that this notion of sample complexity in data attribution is a commonly used efficiency metric in the field (see e.g, Ilyas et al., (2022); Park et al., (2023)). We emphasize this metric as each retraining constitutes a significant computational effort, potentially scaling with the dataset size NN itself. Hence, it captures the primary bottleneck for the Verifier. We are now ready to state the theorem.
@@ 3 A “Simple” Non-Interactive Verification Protocol
While this cost is independent of the dataset size NN, the cubic dependence on $1/\varepsilon$ can still be substantial.
@@ 4.1 Proof Sketch of Main Theorem
Efficiency. To show efficiency, we will directly estimate the expected number of model trainings done by the Verifier, which is straightforward: it is the sum of the number of model trainings in Round 33, which consists of $O(\log(1/\delta)/\epsilon^{2})$ spot checks, and $O(\log(1/\delta)/\epsilon^{2})$ to estimate the MSE. Thus the overall cost is $O(\log(1/\delta)/\epsilon^{2})$.
@@ A.1 Harmonic Analysis on the p-biased Cube
We consider the $p\text{-biased}$ distribution $\mathcal{B}_{p}$ over $\{\pm 1\}^{N}$, where each coordinate $x_{i}$ is independently chosen to be 11 with probability $p$ and 1-1 with probability 1p1-p. Let $\mu=\mathbb{E}\left[x_{i}\right]=p-(1-p)=2p-1$ and $\sigma^{2}=\mathrm{Var}(x_{i})=\mathbb{E}\left[x_{i}^{2}\right]-(\mathbb{E}\left[x_{i}\right])^{2}=1-(2p-1)^{2}=4p(1-p)$. The uniform distribution corresponds to $p=1/2$, where $\mu=0$ and $\sigma^{2}=1$.
@@ Appendix B Efficient Estimation of Linear Datamodel Residual
This approach allows estimating the quality of the best linear fit using a number of samples that is proportional to $1/\varepsilon^{3}$ for a desired approximation error $\varepsilon$ but is independent of the dimension NN.